Page Two


The RF/IF chassis is removed and here is the underneath. The view is quite similar to on the earlier TV22 but this version uses 9-pin EF80 valves rather than 7-pin EF91's.


A preview of where much of the future work will take place, with this underneath view of the main chassis.


The chassis parts now receive their preliminary wash in a detergent solution, to remove the worst of the dirt. This is helped along with a stiff brush - and WD40 when there are waxy deposits. I have recently got Parkinson's Disease and can now call upon its shaking feature for more intense brushing!

There are no exceptions... the line output transformer and other coils and chokes are washed too. However, all are thoroughly blown over with compressed air straight afterwards.

Here we see the focus magnet being dunked...


I've been thinking more about the eventual use of this set. It would be nice if I could once again view my collection of old 405-line and 625-line VHS cassettes. There's some priceless archive materials amongst them. There's another row of them hidden behind this lot.

So I've decided to include baseband 405-line and 625-line inputs as well as the normal Channel 1 405. The set will be switchable between the three modes. Stay tuned to see how this is done...


During cleaning and examination of the LOPT, I discover the whisker of wire that runs from the EHT overwind to the EHT rectifier has come adrift.

Now this may be my fault, due to carelessness. Anyway the rule is "if you broke it you fix it". There is now nothing visible left of the end of the winding, but after melting the wax with a soldering iron and some gentle prodding around I think I can see a tiny wire hair about 1mm long that represents what I am looking for.

I pass a new connecting wire through the EHT rectifier's mounting hole and curl it over the top of the board to discourage it from moving during soldering. Then, ( 'X' marks just below the spot) the new wire is bent to the exact position and solder applied.

At the second attempt, it works! I can see the correct 350-ohm reading across the overwind. But only actual use will prove whether this repair proves durable.


Here's the LOPT with the rotten valve top cap feed wires replaced and a new EY51 EHT rectifier fitted.

Elsewhere, I've replaced the perished wires feeding the frame scan coils with new rubber wires of the correct colours.

All the little sub-assemblies that were removed from the chassis are now serviced.


The perished remains of the speaker leads, seen at the left side of the transformer, are now pulled off.

The purple covered main carrying lead, to the upper left, is in dangerous condition and will be replaced in due course.


I now carry out a survey of all the capacitor values and quantities that will be needed to complete the job. Here is the list being compared in stores. It turns out I have enough of everything in stock...


One capacitor harder to source would have been the main (double) HT smoothing electrolytic. However both parts of this prove susceptible to re-forming. This is carried out overnight.

Here we see, toward the end of the process, the higher capacitance (250uF) section showing a leakage current is down to 181 uA. This figure is very good.


Next job, to re-cap the RF/IF unit...

The pot toward the middle is the Vision Interference Limiter. In practice, this won't be used but nevertheless it is pumped full of Waxoyl to prevent any future scratchiness.


But what's this? Someone has snipped this capacitor connection in the past.

This is actually one of the heater chain decoupling capacitors and doesn't do very much. It will be replaced later anyway.


All the waxies underneath are duly replaced. This is to be a set that will be in regular use and I don't want it going wrong regularly as the price for maintaining strict originality.

Dirt is removed from crannies with a Q-tip impregnated with WD40. The valve bases are also sprayed with WD40 before wiping over. This will have a temporary effect on their electrical characteristics but they will have dried out and be back to normal long before the set is switched on.


Around now, all of the many pots on the chassis are injected with waxoyl, for a scratch-free future.

The final heater chain dropper lives underneath on this set and must have been gently smoking at one time, since there is a black deposit on the chassis above, all round it. The dropper is temporarily removed - here's the mounting hole - and the deposit cleaned off.


This is the original frame output cathode decoupling capacitor. It is returned to +HT, not 'deck'. I am not sure why this is, though doing it this way will tend to cancel any HT hum on the frame scan. It is replaced.

Note the 'Oct 53' date on its case. This set would not have been in the shops in time for the Coronation!


Next this gubbins must come off! This is the back view of the mains adjustment dropper assembly. The rubber wiring will be mostly replaced.


Getting it off involves removal of some other parts underneath to give access to the mounting screws. In general the purple mains-carrying rubber insulated wiring is in perished condition and will have to go.